Problem 19.44 A +5.0-uC charge experiences a o.56-N force in the positive y direction. Part A If this charge is replaced with a-2.7-C charge, find the magnitude of the force in this case. Express your answer using two significant figures. ANSWER: 1/22/2019 Chapter 19 Part B Find the direction of the force in this case. ANSWER O in the positive y direction O in the negative y direction O perpendicular to the y direction
Answer #1

given charge Q= 5,0μέ = 5.010c

force F= 0.56 N

we know F = Q* E ( where E= electric field)............(1)

so E= F/Q

= 0.56M)/5.0 * 10-6-(0.112 * 106N/c) ................(2)

now the charge is replaced by (Q') = -2.7mu c = 2.7 10c

using equation .........(1) and substituting the value of E from equation........(2) we can find the new force.

a) So F = Q'E = (-2.7 * 10-С) * (0.112 * 10°M/C)

F0.3024Ny

b) as we have negative sign in force so the force is in the negative y direction.